The Rockford Boat Company just declared a dividend of $4 per share. It is expected that the company's earnings and dividends will grow at a rate of 20% next year, 15% the following year and 11% the year after that. After that, earnings and dividends are expected to grow at a constant rate of 3%. Required: If you expect a rate of return of 10% on this stock, what would you expect to pay for a share of Rockford Boat Company stock today?
Answer #1
Fair market value of the stock today is the PV of the expected dividends
when discounted at the required rate of return of 10%.
Year Dividend PVIF at 10% PV at 10%
0 $           4.000
1 $           4.800 0.90909 $                4.36
2 $           5.520 0.82645 $                4.56
3 $           6.127 0.75131 $                4.60
Sum of PV of dividends of years 1 to 3 $              13.53
Continuing value of dividends at t3 = 6.127/(0.10-0.03) = $              87.53
PV of continuing value = 87.53*0.75131 = $              65.76
Market value of the stock today = 13.53+65.76 = $              79.29